Deb is 4 years younger than Beth. Six years ago the sum of their ages was 46. How old are they now? I'm unfaliar with finding answer when asked "six years ago." How do I set this up? Thanks.
Answer by stanbon(75887)

Deb is 4 years younger than Beth. Six years ago the sum of their ages was 46. How old are they now? I'm unfamiliar with finding answer when asked "six years ago."
------------
Now Eq: D = B - 4
6yrs ago Eq::: D-6 + B-6 = 46
--------------------
Simplify:
D - B = -4
D + B = 58
----
2D = 54
Deb = 27 yrs (Deb's age now)
B = 27-4 = 23 yrs (Beth's age now)
===========
Cheers,
Stan H.